OpenStudy (anonymous):

Every day, there are 4 times more likes on an internet video of a horse which is modeled by the function c(n) = (4)n − 1, where n is the number of days since the video posted. On the first day, there were 100 likes. What is the function that shows the number of likes each day?

OpenStudy (anonymous):

@beccaboo333

OpenStudy (beccaboo333):

Oh I'm not good with math

OpenStudy (anonymous):

oh ok do you know anyone that is that's online?

OpenStudy (beccaboo333):

@amistre64 is good :3

OpenStudy (amistre64):

lol, im good?

OpenStudy (anonymous):

:D ill be the judge of that lol jkjkjk

OpenStudy (beccaboo333):

Amistre is the bestest person with math in the whole world. Trust what he says to do for math. Better?

OpenStudy (amistre64):

so, how many likes on the first day?

OpenStudy (amistre64):

each day is 4 times more than c(n) = (4)n − 1 .... you are going to have to notate this better.

OpenStudy (anonymous):

c(n)=4^( n-1) sorry

OpenStudy (amistre64):

Every day, there are 4 times more likes than c(n) = 4^(n-1), where n is the number of days since the video posted. On the first day, there were 100 likes. On our own video? or on the video of the horse? What is the function that shows the number of likes each day?

OpenStudy (anonymous):

so is it 100(4)^n-1?

OpenStudy (amistre64):

c1 = 1 c2 = 4 c3 = 16 c4 = 64 etc is the number of likes on the horse video, not ours ... right?

OpenStudy (anonymous):

yes :)

OpenStudy (amistre64):

and we are 4 times more than c(n) each day.

OpenStudy (amistre64):

but we start off our first day with 100 likes .... the informations not making any sense to me since 100 on day 1 is clearly not 4*1

OpenStudy (anonymous):

yah

OpenStudy (amistre64):

What is the function that shows the number of likes each day? Given that: Every day, there are 4 times more likes on an internet video of a horse which is modeled by the function c(n) = 4^(n − 1), where n is the number of days since the video posted. and: On the first day, there were 100 likes. it appears to say that the number of likes each day is 4c(n) since the function c(n) can already be assessed. But then it suggests that that is not the case ....

OpenStudy (amistre64):

i cant make any good sense of it

OpenStudy (anonymous):

poop haha what would you do? the answer choices are c(n) = 100(4)^n − 1 c(n) = (4)(100)^(n − 1) c(n) = (100)^n − 1 c(n) = (4)^100 − 1

OpenStudy (amistre64):

id take a wild guess .... the first one

OpenStudy (anonymous):

and I have a couple more can you help?

OpenStudy (amistre64):

not if they are anything like this one

OpenStudy (anonymous):

hahaha umm

OpenStudy (amistre64):

oh good, numbers .... thats much better

OpenStudy (amistre64):

i can narrow it down to 1, which when chked fits. any ideas?

OpenStudy (amistre64):

notice the graph decreases as n gets bigger. 16^n doesnt decrease 2^n doesnt decrease (-2)^n doesnt decrease (1/2)^n does decrease

OpenStudy (anonymous):

1/2^n-1 :D

OpenStudy (anonymous):

@amistre64 :D